### v1

@@ 3.2.2 Neural operators
for $h\in\mathcal{H}$ and $x\in\Omega$. The particular choice of parametrization of the matrix-valued kernel $\kappa_{t}\colon\Omega\times\Omega\to\mathbb{R}^{d_{\mathrm{c}}\times d_{\mathrm{c}}}$ determines the classification of the neural operator architecture. For example, choosing $\kappa_{t}$ to be a neural network localized to patches defines the Graph Kernel Neural Operator, while choosing a low-rank kernel $\kappa_{t}$ with learnable factors leads to the Low-Rank Neural Operator [147, Sec. 4]. There are generalizations to nonlinear $\mathcal{K}_{t}$ in which its integral kernel $\kappa_{t}$ depends on the input function $h$ itself [147]. Regardless of the parametrization used, what is important is that $\mathcal{K}_{t}$ is nonlocal [172].
This chapter emphasizes the use of the Fourier Neural Operator (FNO) [24, 189], which corresponds to a neural operator $\Psi_{\mathrm{FNO}}$ of the form (3.11) on the unit hypertorus $\Omega=\mathbb{T}^{d}\simeq[0,1]^{d}_{\mathrm{per}}$ with integral operator parametrization
@@
This corresponds to enforcing that the matrix-valued function $\kappa_{t}$ in (3.15) is a translation-invariant convolution kernel parametrized directly in Fourier space by its Fourier series coefficients $P^{(k)}_{t}\in\mathbb{C}^{d_{c}\times d_{c}}$ for each $k\in\mathbb{Z}^{d}$.
A practical discretization of the FNO is usually performed on equally-spaced grids with the Fast Fourier Transform algorithm after truncating the series in (3.16) to a maximum number of resolved Fourier modes [190, 189]. It is also possible to apply the FNO to functions on non-periodic domains [191, 78, 24, 25, 172, 192]. In addition to its vast practical success in scientific applications, theory for the FNO is also beginning to emerge [192, 190, 183, 172, 78].
@@ 4 Learning inverse problem solvers
However, a direct, off-the-shelf application of operator learning to inverse problems encounters several immediate challenges. Beyond the need for a sufficiently rich dataset to support the costly offline training step (3.2), the forward map $\mathcal{G}$ is typically neither invertible nor well-conditioned. As a result, one cannot simply define the target operator to be $\mathsf{F}=\mathcal{G}^{-1}$ without carefully restricting the domain of $\mathcal{G}$ to a sufficiently small set in order for a continuous inverse to be well-defined. Ensuring that the data distribution assigns probability one to such a set is itself a nontrivial challenge. Moreover, the additive noise model (3.1) commonly assumed in forward operator learning is inappropriate for the inverse setting. Here, the roles of input and output must be reversed: the inverse map is trained to recover noise-free ground-truth parameters, while the noisy outputs of the forward map now serve as inputs. This reversal leads to a challenging errors-in-variables learning problem [193, 194, 195].
Despite these challenges, there is growing theoretical and empirical evidence that estimating the inverse map with operator learning is both statistically principled and practically robust [173, 196, 25, 197, 198]. Neural operators can be trained using loss functions that explicitly target desired solution attributes, such as accurate parameter recovery or physical consistency with the forward model. In contrast, plug-in estimators that invert a learned forward map are not goal-oriented and therefore cannot achieve optimal performance; see [196, Eqn. (219), p. 58] for details. Moreover, even when the underlying inverse problem is severely ill-posed, the average-case nature of the training process tends to suppress the impact of worst-case instability: the extreme inputs and outputs responsible for instability are unlikely under realistic data distributions. End-to-end inverse map learning is especially advantageous in scenarios where data is plentiful and the forward or measurement processes are imperfect or only partially known.
A subset of the research on learning inverse solvers distinguishes itself by blending knowledge of the forward operator into the design of algorithms [25, 197, 199, 200, 6, 7, 8, 17, 201], as advocated by [202]. This could involve correcting errors in the forward model or inverse operator with learned postprocessing maps [26, 49]. Other work develops invertible network architectures that simultaneously learn approximate forward and inverse maps while incorporating physics information [203, 204, 205, 206, 9, 10, 207]. It is also possible to adopt a distribution-matching approach for the inverse map loss component [208], which lifts similar ideas [209, 100, 102, 101] to the operator learning setting.
@@ 4.1 Mathematical formulation
This subsection aims to make the inverse map learning problem more precise. We work under the model (2.1), that is, $y=\mathcal{G}(u)+\eta$. Suppose that $(\mathcal{U},\mathsf{d}_{\mathcal{U}})$ is a metric space of parameters and $(\mathcal{Y},\mathsf{d}_{\mathcal{Y}})$ is a metric space of data. Both spaces are assumed to be infinite-dimensional. We view
@@
as a mapping from its domain $(\operatorname{Dom}(\mathcal{G}),\mathsf{d}_{\mathcal{U}})$ into $\mathcal{Y}$. The introduction of the set $\operatorname{Dom}(\mathcal{G})\subseteq\mathcal{U}$ is necessary in many inverse problems because although $\mathcal{G}$ may be well-defined on the whole of the metric space $\mathcal{U}$, often it is only injective on a smaller set (thus guaranteeing unique solutions of the noiseless inverse problem). This is the case for EIT from Example 1, where even smaller sets are identified such that the restriction of $\mathcal{G}$ to these sets has a uniformly continuous inverse [210]. Thus, this subsection assumes that $\operatorname{Dom}(\mathcal{G})$ is chosen such that the inverse map $\Psi^{\star}$ defined by
@@
The goal of end-to-end inverse map learning is to approximate the operator $\Psi^{\star}$ by some $\widehat{\Psi}$ in the sense that $\widehat{\Psi}(\mathcal{G}(u)+\eta))\approx u$ for $(u,\eta)$ ranging over an appropriate set. An operator learning framework that seeks to achieve this goal is presented in Section 4.1.1. It is important to appreciate that the true inverse map $\Psi^{\star}$ in (4.2) is only defined on the range of $\mathcal{G}$. Section 4.1.2 explores the implications of this fact both in theory and in practice.
@@ 4.1.1 Training and evaluating learned inverse maps
is a natural notion of accuracy. However, we cannot expect (4.6) to be small unless the noise $\eta$ is small in an appropriate sense. Other choices include the noise-free parameter reconstruction loss
@@
or the data fitting losses
@@ 4.1.2 Noisy data and the range of the forward map
Conceptually, an ideal approximation $\Psi$ of $\Psi^{\star}$ would take the form $\Psi\coloneqq\Psi^{\star}\circ P_{\mathcal{G}}$, where $P_{\mathcal{G}}\colon\mathcal{Y}\to\operatorname{Ran}(\mathcal{G})$ is some notion of projection onto $\operatorname{Ran}(\mathcal{G})$. In this way, $\Psi\colon\mathcal{Y}\to\mathcal{U}$ extends $\Psi^{\star}$: $\Psi(y)=\Psi^{\star}(y)$ if $y\in\operatorname{Ran}(\mathcal{G})$. If $y\notin\operatorname{Ran}(\mathcal{G})$, then $P_{\mathcal{G}}(y)$ is a “denoised” version of $y$ in a sense. However, it is practically problematic to construct $P_{\mathcal{G}}$ because for many infinite-dimensional nonlinear inverse problems, $\operatorname{Ran}(\mathcal{G})$ is typically a complicated nonconvex set that is difficult to characterize [58]. Nonetheless, it is intuitively possible that by minimizing the end-to-end ERM training objective (4.5), $\widehat{\Psi}$ is implicitly approximating the extension $\Psi$ even though $\Psi$ is not accessible. An interesting research direction is to learn neural operators that explicitly project onto $\operatorname{Ran}(\mathcal{G})$ by training on enough data from the range.
@@
The first term in the upper bound (4.9) is controlled by how well $\widehat{\Psi}$ approximates $\widetilde{\Psi}^{\star}$. The second error term is the modulus of continuity of the learned inverse map evaluated at the maximal noise level; this term tends to zero as $\delta\to 0$. While it is a hard problem to quantitatively control Lipschitz or other uniform continuity properties of trained neural networks and neural operators, there do exist specialized architectures that attempt this program [216, 198].
@@
One way to decouple the role of regularization from inverse map training is to build regularization into the training data itself [217]. This method trains approximate inverse maps on data consisting of regularized inverse problem solutions. These data are obtained from classical inverse solvers, e.g., Tikhonov or Bayesian regularization. The insight is that training on pre-regularized data corresponds to learning a regularized inverse map instead of $\Psi^{\star}$ [217, 197]. While this addresses the noisy data problem, the target regularized inverse no longer exactly solves the inverse problem on the range of $\mathcal{G}$. That is, it does not extend the true inverse. This reflects a tradeoff between accuracy and regularization in the operator learning setting. Similar considerations arise when choosing to define the inverse map outside of its domain as a point estimator of a Bayesian posterior, such as the posterior mean [218].
@@ 4.2.1 Fourier Neural Mappings
for each input vector $z\in\mathbb{R}^{d_{\mathrm{c}}}$ and point $x\in\Omega$ in the domain. In (4.11), $\kappa\colon\Omega\to\mathbb{R}^{d_{\mathrm{c}}\times d_{\mathrm{c}}}$ is a learnable matrix-valued function. Thus, $\mathscr{D}z\colon\Omega\to\mathbb{R}^{d_{\mathrm{c}}}$ is a vector-valued function. In particular, Fourier Neural Mappings (FNMs) generalize FNOs—which only map functions to functions—by accommodating VtF or function-to-vector maps. In the former case, these Fourier Neural Decoders parametrize the function $\kappa$ in Fourier space by its conjugate symmetric Fourier coefficients $\{P^{(k)}\}_{k\in\mathbb{Z}^{d}}\subseteq\mathbb{C}^{d_{\mathrm{c}}\times d_{\mathrm{c}}}$. Letting $\Omega=\mathbb{T}^{d}$, this leads to the decoder
@@
where $\mathscr{D}$ is given in (4.12) and $S\colon\mathbb{R}^{J}\to\mathbb{R}^{d_{\mathrm{c}}}$ is a shallow neural network; compare to (3.11), which uses $\mathscr{L}_{1}$ instead of $\mathscr{D}$ and $\mathcal{S}$ instead of $S$. The $\{\mathscr{L}_{t}\}$ are the usual Fourier integral operator layers as in (3.14) and (3.16). The map $\mathcal{Q}$ is as in (3.13). FNMs retain the computational benefits and universal approximation properties of FNOs [78, Sec. 3].
If the measurements in the inverse problem are comprised of both functions and finite vectors, then the extension of FNMs introduced in [219] can be used to handle such multimodal data. The function-to-vector ideas from FNMs [78] are also useful when the dimension of the input and output spatial domains do not match, such as in inverse boundary value problems (IBVPs) [211]. Finally, we note that if $J$ is allowed to vary, it could be more natural to model the inverse operator as a function-to-function map (Section 3.2), operator-to-function map (Section 4.2.2), or measure-to-function map (Section 4.2.3), depending on the nature of the limiting object as $J\to\infty$. This removes the need to re-train the model every time $J$ changes, which would be the case for VtF architectures.
@@ 4.2.2 Operator Recurrent Neural Networks
for each $j\in\{0,1\}$ and every $h$, $\Lambda$, $t$, and $x$. The $\{\mathcal{K}_{t}^{(j)}\}$ and $\{\widetilde{\mathcal{K}}_{t}^{(j)}\}$ are nonlocal linear kernel integral operators of the form (3.15) or (3.16). Notice how $\Lambda$ is introduced in the previous display via operator composition, i.e., matrix multiplication. If the activation functions $\{\sigma_{t}\}$ are set to the identity, then a deep ORN-Net computes operator polynomials and starts to resemble a truncated Neumann series expansion [196, Sec. 6]. Thus, the nonlinear version with non-polynomial activation functions generalizes the network’s representation capacity to more complex operator functions.
Extensions to the core architecture include adding memory to the layers instead of settling for the Markovian evolution (4.15); see [196, Sec. 2.2–2.3]. The architecture design is further motivated by specific inverse problem structure such as the compositional or iterative nature of the ground truth inverse problem solution map. This leads to inductive biases that promote compactness, sparsity, and low-rankness in the model, which in turn lead to improved statistical generalization bounds. We refer to [196, Sec. 2–3, 7] for the details and an application to boundary control of the wave equation. We note that both the quantitative universal approximation and statistical learning theory for ORN-Net [196] only handles target functions mapping $\mathbb{R}^{n\times n}$ to $\mathbb{R}^{m}$ and does not carry over immediately to the infinite-dimensional case. This is an important avenue for future mathematical research. Numerical implementation of ORN-Nets in both finite- and infinite-dimensional settings is also an important future direction to determine the practical benefits of the architecture’s multiplicative structure.

### v2

@@ 3.2.2 Neural operators
for $h\in\mathcal{H}$ and $x\in\Omega$. The particular choice of parametrization of the matrix-valued kernel $\kappa_{t}\colon\Omega\times\Omega\to\mathbb{R}^{d_{\mathrm{c}}\times d_{\mathrm{c}}}$ determines the classification of the neural operator architecture. For example, choosing $\kappa_{t}$ to be a neural network localized to patches defines the Graph Kernel Neural Operator, while choosing a low-rank kernel $\kappa_{t}$ with learnable factors leads to the Low-Rank Neural Operator [157, Sec. 4]. Alternatively, expanding $\kappa_{t}$ in a wavelet basis or Laplace–Beltrami operator eigenfunctions on a manifold leads to the Wavelet Neural Operator [203, 204] and the Neural Operator on Riemannian Manifolds [205], respectively. There are generalizations to nonlinear $\mathcal{K}_{t}$ in which its integral kernel $\kappa_{t}$ depends on the input function $h$ itself [157]. Regardless of the parametrization used, what is important is that $\mathcal{K}_{t}$ is nonlocal [182].
In the context of inverse problems, this chapter later makes use of the Fourier Neural Operator (FNO) [26, 206], which corresponds to a neural operator $\Psi_{\mathrm{FNO}}$ of the form (3.13) on the unit hypertorus $\Omega=\mathbb{T}^{d}\simeq[0,1]^{d}_{\mathrm{per}}$ with integral operator parametrization
@@
This corresponds to enforcing that the matrix-valued function $\kappa_{t}$ in (3.17) is a translation-invariant convolution kernel parametrized directly in Fourier space by its Fourier series coefficients $P^{(k)}_{t}\in\mathbb{C}^{d_{c}\times d_{c}}$ for each $k\in\mathbb{Z}^{d}$. A practical discretization of the FNO is usually performed on equally-spaced grids with the Fast Fourier Transform algorithm after truncating the series in (3.18) to a maximum number of resolved Fourier modes [207, 206]. While this limits the most popular implementations to simple periodic box domains, recent works have extended the FNO to non-periodic domains and complex geometries [59, 208, 82, 26, 27, 182, 209, 210]. Another limitation includes the FNO’s handling of spatially local phenomena [211, 212]. Despite these particular drawbacks that deserve future attention, the FNO and its variants have demonstrated vast practical success in scientific applications. Theory for the FNO is also beginning to emerge [209, 207, 193, 182, 82].
@@ 4 Learning inverse problem solvers
However, a direct, off-the-shelf application of operator learning to inverse problems encounters several immediate challenges. Beyond the need for a sufficiently rich dataset to support the costly offline training step (3.2), the forward map $\mathcal{G}$ may not be invertible, or if it is, its inverse may be ill-conditioned. As a result, one cannot simply define the target operator to be $\mathsf{F}=\mathcal{G}^{-1}$ without carefully restricting the domain of $\mathcal{G}$ to a sufficiently small set in order for a continuous inverse to be well-defined. Ensuring that the data distribution assigns probability one to such a set is itself a nontrivial challenge. Moreover, the additive noise model (3.1) commonly assumed in forward operator learning is inappropriate for the inverse setting. Here, the roles of input and output must be reversed: the inverse map is trained to recover noise-free ground-truth parameters, while the noisy outputs of the forward map now serve as inputs. This reversal leads to a challenging errors-in-variables learning problem [213, 214, 215].
Despite these challenges, there is growing theoretical and empirical evidence that estimating the inverse map with operator learning is both statistically principled and practically robust [183, 216, 27, 217, 218]. Neural operators can be trained using loss functions that explicitly target desired solution attributes, such as accurate parameter recovery or physical consistency with the forward model. In contrast, plug-in estimators that invert a learned forward map are not goal-oriented and therefore cannot achieve optimal performance; see [216, Eqn. (219), p. 58] for details. Moreover, even when the underlying inverse problem is severely ill-posed, the average-case nature of the training process tends to suppress the impact of worst-case instability: the extreme inputs and outputs responsible for instability are unlikely under realistic data distributions. End-to-end inverse map learning is especially advantageous in scenarios where data is plentiful and the forward or measurement processes are imperfect or only partially known.
A subset of the research on learning inverse solvers distinguishes itself by blending knowledge of the forward operator into the design of algorithms [27, 217, 219, 220, 6, 7, 8, 17, 221], as advocated by [222]. This could involve correcting errors in the forward model or inverse operator with learned postprocessing maps [28, 51]. Other work, such as [223, 224, 225], develops injective architectures that account for the stability of the inverse problem. Additional approaches introduce invertible networks that jointly learn approximate forward and inverse maps while incorporating physical information [226, 227, 228, 9, 10, 229]. It is also possible to adopt a distribution-matching approach for the inverse map loss component [230], which lifts similar ideas [231, 106, 108, 107] to the operator learning setting.
@@ 4.1 Mathematical formulation
This subsection aims to make the inverse map learning problem more precise. We work under the additive model (2.1), that is, $y=\mathcal{G}(u)+\eta$. Suppose that $\mathcal{U}$ is a Banach space of parameters and $\mathcal{Y}$ is a Banach space of data. Let $\mathsf{d}_{\mathcal{U}}$ and $\mathsf{d}_{\mathcal{Y}}$ denote the metrics induced by the norms on $\mathcal{U}$ and $\mathcal{Y}$, respectively. Both $\mathcal{U}$ and $\mathcal{Y}$ are assumed to be infinite-dimensional. We view
@@
as a mapping from its domain $(\operatorname{Dom}(\mathcal{G}),\mathsf{d}_{\mathcal{U}})$ into $\mathcal{Y}$. The introduction of the set $\operatorname{Dom}(\mathcal{G})\subseteq\mathcal{U}$ is necessary in many inverse problems because although $\mathcal{G}$ may be well-defined on the whole of the Banach space $\mathcal{U}$, often it is only injective on a smaller set (thus guaranteeing unique solutions of the noiseless inverse problem). This is the case for EIT from Example 1, where even smaller sets are identified such that the restriction of $\mathcal{G}$ to these sets has a uniformly continuous inverse [232]. Thus, this subsection assumes that $\operatorname{Dom}(\mathcal{G})$ is chosen such that the inverse map $\Psi^{\star}$ defined by
@@
The goal of end-to-end inverse map learning is to approximate the operator $\Psi^{\star}$ by some $\widehat{\Psi}$ in the sense that $\widehat{\Psi}(\mathcal{G}(u)+\eta)\approx u$ for $(u,\eta)$ ranging over an appropriate set. An operator learning framework that seeks to achieve this goal is presented in Section 4.1.1. It is important to appreciate that the true inverse map $\Psi^{\star}$ in (4.2) is only defined on the range of $\mathcal{G}$. Section 4.1.2 explores the implications of this fact both in theory and in practice.
@@ 4.1.1 Training and evaluating learned inverse maps
is a natural notion of accuracy. However, we cannot expect (4.6) to be small unless the noise $\eta$ is small in an appropriate sense. In statistical learning, this is addressed by instead considering the excess risk [158, Sec. 2.3]. Other choices include the noise-free average parameter reconstruction loss
@@
or the average data fitting losses
@@ 4.1.2 Noisy data and the range of the forward map
Conceptually, a natural approximation $\Psi$—which in some cases will be optimal—of $\Psi^{\star}$ would take the form $\Psi\coloneqq\Psi^{\star}\circ P_{\mathcal{G}}$, where $P_{\mathcal{G}}\colon\mathcal{Y}\to\operatorname{Ran}(\mathcal{G})$ is some notion of projection onto $\operatorname{Ran}(\mathcal{G})$. In this way, $\Psi\colon\mathcal{Y}\to\mathcal{U}$ extends $\Psi^{\star}$, i.e., $\Psi(y)=\Psi^{\star}(y)$ if $y\in\operatorname{Ran}(\mathcal{G})$. If $y\notin\operatorname{Ran}(\mathcal{G})$, then $P_{\mathcal{G}}(y)$ is a “denoised” version of $y$ in a sense. However, it is practically problematic to construct $P_{\mathcal{G}}$ because for many infinite-dimensional nonlinear inverse problems, $\operatorname{Ran}(\mathcal{G})$ is typically a complicated nonconvex set that is difficult to characterize [62]. Nonetheless, it is intuitively possible that by minimizing the end-to-end ERM training objective (4.5), $\widehat{\Psi}$ is implicitly approximating the extension $\Psi$ even though $\Psi$ is not accessible. An interesting research direction is to learn neural operators that explicitly project onto $\operatorname{Ran}(\mathcal{G})$ by training on enough data from the range.
@@
The first term in the upper bound (4.9) is controlled by how well $\widehat{\Psi}$ approximates $\widetilde{\Psi}^{\star}$. The second error term is the modulus of continuity of the learned inverse map evaluated at the maximal noise level; this term tends to zero as $\delta\to 0$. While it is a hard problem to quantitatively control Lipschitz or other uniform continuity properties of trained neural networks and neural operators, there do exist specialized architectures that attempt this program [237, 218, 238, 239].
@@
One way to decouple the role of regularization from inverse map training is to build regularization into the training data itself [240]. This method trains approximate inverse maps on data consisting of regularized inverse problem solutions. These data are obtained from classical inverse solvers, e.g., Tikhonov or Bayesian regularization. The insight is that training on pre-regularized data corresponds to learning a regularized inverse map instead of $\Psi^{\star}$ [240, 217]. While this addresses the noisy data problem, the target regularized inverse no longer exactly solves the inverse problem on the range of $\mathcal{G}$. That is, it does not extend the true inverse. This reflects a tradeoff between accuracy and regularization in the operator learning setting. Similar considerations arise when choosing to define the inverse map outside of its domain as a point estimator of a Bayesian posterior, such as the posterior mean [241]. In practice, trained inverse maps will not extend the true inverse map, regardless of whether the training data is pre-regularized or not.
@@ 4.2.1 Fourier Neural Mappings
for each input vector $z\in\mathbb{R}^{d_{\mathrm{c}}}$ and point $x\in\Omega$ in the domain. In (4.11), $\kappa\colon\Omega\to\mathbb{R}^{d_{\mathrm{c}}\times d_{\mathrm{c}}}$ is a learnable matrix-valued function. Thus, $\mathscr{D}z\colon\Omega\to\mathbb{R}^{d_{\mathrm{c}}}$ is a vector-valued function. In particular, Fourier Neural Mappings (FNMs) generalize FNOs—which only map functions to functions—by accommodating VtF or function-to-vector (FtV) maps. In the former case, these Fourier Neural Decoders parametrize the function $\kappa$ in Fourier space by its conjugate symmetric Fourier coefficients $\{P^{(k)}\}_{k\in\mathbb{Z}^{d}}\subseteq\mathbb{C}^{d_{\mathrm{c}}\times d_{\mathrm{c}}}$. Letting $\Omega=\mathbb{T}^{d}$, this leads to the decoder
@@
where $\mathscr{D}$ is given in (4.12) and $S\colon\mathbb{R}^{J}\to\mathbb{R}^{d_{\mathrm{c}}}$ is a shallow neural network; compare to (3.13), which uses $\mathscr{L}_{1}$ instead of $\mathscr{D}$ and $\mathcal{S}$ instead of $S$. The $\{\mathscr{L}_{t}\}$ are the usual Fourier integral operator layers as in (3.16) and (3.18). The map $\mathcal{Q}$ is as in (3.15). FNMs retain the computational benefits and universal approximation properties of FNOs [82, Sec. 3].
If the measurements in the inverse problem are comprised of both functions and finite vectors, then the extension of FNMs introduced in [242] can be used to handle such multimodal data. The FtV ideas from FNMs [82] are also useful when the dimension of the input and output spatial domains do not match, such as in inverse boundary value problems (IBVPs) [59]. Finally, we note that if $J$ is allowed to vary, it could be more natural to model the inverse operator as a function-to-function map (Section 3.2), operator-to-function map (Section 4.2.2), or measure-to-function map (Section 4.2.3), depending on the nature of the limiting object as $J\to\infty$. This removes the need to re-train the model every time $J$ changes, which would be the case for VtF architectures.
@@ 4.2.2 Operator Recurrent Neural Networks
for each $j\in\{0,1\}$ and every $h$, $\Lambda$, $t$, and $x$. The $\{\mathcal{K}_{t}^{(j)}\}$ and $\{\widetilde{\mathcal{K}}_{t}^{(j)}\}$ are nonlocal linear kernel integral operators of the form (3.17) or (3.18). Notice how $\Lambda$ is introduced in the previous display via operator composition, i.e., matrix multiplication. If the activation functions $\{\sigma_{t}\}$ are set to the identity, then a deep ORN-Net computes operator polynomials and starts to resemble a truncated Neumann series expansion [216, Sec. 6]. Thus, the nonlinear version with non-polynomial activation functions generalizes the network’s representation capacity to more complicated operator functions.
Extensions to the core architecture include adding memory to the layers instead of settling for the Markovian evolution (4.15); see [216, Sec. 2.2–2.3]. The architecture design is further motivated by specific inverse problem structure such as the compositional or iterative nature of the ground truth inverse problem solution map. This leads to inductive biases that promote compactness, sparsity, and low-rankness in the model, which in turn lead to improved statistical generalization bounds. We refer to [216, Sec. 2–3, 7] for the details and an application to boundary control of the wave equation. We note that both the quantitative universal approximation and statistical learning theory for ORN-Net [216] only handles target functions mapping $\mathbb{R}^{n\times n}$ to $\mathbb{R}^{m}$ and does not carry over immediately to the infinite-dimensional case. This is an important avenue for future mathematical research. Numerical implementation of ORN-Nets in both finite- and infinite-dimensional settings is also an important future direction to determine the practical benefits of the architecture’s multiplicative structure.